### v1

# The central engine of low-luminosity radio galaxy 3C 270

Abstract
We present the polarization spectra of the nucleus of 3C 270. We confirm that the polarization angle of both the continuum and the emission lines are close to perpendicular to the jet direction after careful correction of interstellar polarization, which indicates polar scattering. The Stokes flux spectrum resembles the total flux spectrum, with no need for a broad component from the Broad Line Region. Over 94% of the Seyfert I type broad line profile would be significantly detected if present in our polarized flux spectrum. We favor that we are mostly likely observing beamed synchrotron instead of Big Blue Bump, and the innermost Narrow Line Region through reflection. This makes 3C 270 the third known case, after NGC 4258 and Centaurus A, where only narrow lines (and beamed continuum, if present) are scattered, with no evidence of an underlying Big Blue Bump.
1 Introduction
The nearby elliptical galaxy NGC 4261 ( = 0.0074) hosts a low-luminosity Active Galactic Nuclei (AGN) and also a bright radio source known as 3C 270. This galaxy is renowned for its prominent dust disc (Jaffe et al., 1993) rotating around a supermassive black hole with a mass of , as reported by Boizelle et al. (2021) and Ruffa et al. (2023).
This famous FR I type radio galaxy has been the subject of numerous studies trying to establish its accretion mode : radiatively efficient, ”thermal” vs. radiatively inefficient, ”non-thermal” (Antonucci, 2012). This interest is motivated by the fact that 3C 270 appears to be a low-luminosity active galaxy showing puzzling evidence for significant accretion power, making it an ideal testing ground for our current understanding of accretion onto supermassive black holes at low Eddington ratios. On one hand, the low-efficiency accretion is supported by emission-line diagnostics from ground-based spectra, which classify 3C 270 as a LINER (Ho et al., 1995). However, HST spectra show Seyfert-type narrow-line ratios on nuclear scales (Ferrarese et al., 1996), indicating the presence of a hard radiation field typical for radiatively efficient accretion.
From Spitzer spectroscopy, Leipski et al. (2009) showed that 3C 270 is dominated by warm thermal dust in the mid-infrared, too warm for stellar-heated dust, indicative of a possible hidden thermal AGN (Ogle et al., 2006). VLT/VISIR data reveal a matching point source, which proves that the very warm thermal dust is indeed of nuclear origin ( pc; Asmus et al., 2011). In some clumpy torus models, a powerful nuclear thermal dust component is not expected for sources of such low luminosity. Instead, the models predict that the torus (and the broad line region) vanishes at low bolometric luminosity ( erg s-1; erg s-1 for 3C 270, Ho, 1999) and/or at low Eddington ratios (Hönig & Beckert, 2007; Elitzur, 2008; Elitzur & Ho, 2009). Additionally, for the low Eddington ratio of 3C 270 (, Ho, 1999), theory predicts advection-dominated accretion (Esin et al., 1997), thus limiting the supply of hard ionizing radiation. However, the nuclear optical emission lines in 3C 270 requires a central engine to produce sufficient ionizing photons. Furthermore, the reprocessed light observed as nuclear warm dust emission necessitates a radiation field sufficient to power the emission. These discrepancies illustrate the need to establish more accurate observational constraints to the models.
One important diagnostic that could place strong constraints on the models for low-luminosity activity is the detection of broad emission lines and Big Blue Bump continuum component. These are generally reliable probes for the presence of a radiatively efficient accretion flow, and the detection of broad lines in polarized light would be the most convincing proof for a hidden thermal AGN. Spectropolarimetry has been used successfully many times to reveal hidden broad lines in narrow-line objects (e.g., Antonucci, 1984; Ogle et al., 1997). Additionally, 3C 270 has been observed spectropolarimetrically by Barth et al. (1999a), surprisingly revealing possibly broadened H in polarized light. Unfortunately, the low signal-to-noise ratio (SNR) precluded a secure detection and width measurement of the broadened line complex. In concert with the detection of warm nuclear dust emission, establishing the presence of broad emission lines in 3C 270 provides strong constraints for our current models of low-luminosity activity, which predict both of these features to be absent. It should be noted however that the theoretical expectations are already know to be contradicted in some cases (e.g., NGC 3147, Bianchi et al., 2019, 2022).
In this work, we present a deeper optical spectropolarimetric observation on 3C 270, aiming to disentangle the H complex in search of hidden broad line. In Section 2, we describe the observation and data reduction and careful removal of ISP. In Section 3, we analyze different models for the polarized flux spectrum. In Section 4, we discuss the implication of accretion mode as from the polarized flux spectrum. The findings are summarized in final Section 5. Throughout the paper, we adopt a CDM cosmology with = 0.7, = 0.3, and = 70 km s-1 Mpc-1.
2 Observation and Data Reduction
2.1 Observation
The observations of 3C 270 were conducted using the FOcal Reducer and low dispersion Spectrograph (FORS2), mounted on the Unit Telescope 1 of the Very Large Telescope (VLT) at the European Southern Observatory (ESO). The observations were spread over three nights: 2014/01/24, 2014/02/28, and 2014/03/01. All data were taken in PMOS mode with the 300V grism.
The observations employed the Half-Wave Plate (HWP) at four different rotation angles (0.0∘, 45.0∘, 90.0∘, and 135.0∘ on the sky plane) in combination with a Wollaston prism to split the light into ordinary (o) and extraordinary (e) rays. Due to the dual-beam design of FORS2, both o-ray and e-ray were recorded simultaneously, with a separation of about 22 arcseconds. The total on-source exposure time for four rotation angles was 1200s, 600s, and 1200s, respectively.
2.2 Data Reduction
The data were reduced using routines written in IDL, following the procedures described by Martínez González et al. (2015). The steps included bias subtraction, flat-field correction, wavelength calibration, sky subtraction, cosmic ray removal, and flux calibration. The o-ray and e-ray were then combined to compute the normalized Stokes parameters and using the following formulas: The normalized Stokes parameters and are defined as:
where is the measured intensity, with the first subscript indicating the ordinary (o) or extraordinary (e) ray, and the second subscript specifying the HWP rotation angle.
The total intensity is computed as: .
The linear polarization degree and polarization angle are derived from and using the standard relations:
For calibration, we selected a zero-polarization standard star (HD 109055) and a polarized standard star (Vela1 95), both observed on 2024/01/24 after the first source observation. Flux and polarimetric standards were part of the default nightly calibration plan.
In the R band, Vela1 95 has a polarization degree of and a polarization angle of 111https://www.eso.org/sci/facilities/paranal/instruments/fors/inst/pola.html. For comparison with the standard, we integrate the normalized Stokes parameters and from our data, weighted by the Johnson R filter transmission. From our measurements, the polarization degree is with a polarization angle of , very close to the standard values. Using the same procedure, we measure the polarization degree of HD 109055, finding , consistent with expected null polarization. The close agreement between our measured and tabulated values indicates that the systemic polarization in the R band is .
The total flux spectrum as well as the Stokes and spectra are shown in Fig.1.
2.3 Interstellar Polarization
Since the continuum polarization is quite low, interstellar polarization (ISP) may significantly impact our interpretation of its origin. Serkowski et al. (1975) demonstrate that the wavelength-dependent polarization degree of the ISP can be accurately modeled using two parameters, and . With the additional constraint of the ISP polarization angle, we can derive the complete ISP information, including the normalized Stokes parameters and . Key parameters for the ISP can be estimated from polarization measurements of stars close to the central source, which provides a local basis for ISP correction.
We start with the stellar polarization catalog by Heiles (2000) and simply take the polarization properties of the closest star in the catalog to the source, while ensuring the star is sufficiently distant so that any dust causing ISP lies between the star and the observer. The best candidate matching these criteria is HD 107289, located pc away from the earth and 0.2∘ from 3C 270 in projection. The measured polarization degree is and polarization angle is . To verify if there is residual dust between HD 107289 and 3C 270 , we use the 3D dust map from Green et al. (2019) to calculate the integrated extinction along the line of sight. In the direction of 3C 270, the integrated extinction is mag, while toward HD 107289, it is mag. This indicates that there is minimal dust beyond HD 107289. Taking , and assuming is 5500 Å, we can then construct the synthetic normalized and spectra as follows:
The computed normalized Stokes parameters are then subtracted from the observed normalized and to derive the intrinsic values. The error of the intrinsic normalized Stokes parameters is calculated by combining the ISP error and the measurement error in quadrature. The normalized Stokes parameter and is in Fig. 2.
3 Results
The radio jet position angle is (Birkinshaw & Davies, 1985). After correcting for the ISP, we find that the polarization angle of the continuum is close to perpendicular to the jet direction, as shown in Fig.3 (see also Table.1). This aligns well with the findings of Capetti et al. (2007), whose broadband polarimetry, tracing primarily continuum polarization, shows a polarization angle of , perpendicular to the jet. Most importantly, the polarization angle of both the H and [S II] complexes, which are less subject to ISP correction, are and , very closely perpendicular to the jet axis at 88∘. Our results are in excellent agreement with theirs, demonstrating the precision of our ISP correction as well.
Capetti et al. (2007) report the optical polarization of a sample of FR I radio galaxies. In particular, their Table 2 shows an overall uncorrelated polarization angles to the jet position angle. The lack of consistent perpendicularity across their sample leads Capetti et al. (2007) to attribute the observed polarization to intrinsic radiation mechanisms, most likely synchrotron emission. However, a notable distinction between 3C 270 and the other FR Is in Capetti et al. (2007), besides its perpendicular polarization, is the high X-ray column density of cm-2 (Zezas et al., 2005; Worrall et al., 2010). Thus, 3C 270 is similar to the FR I prototype Centaurus A, which wasn’t observed by Capetti et al. (2007), but which also has a perpendicular polarization and a high column density associated with the nuclear region, two signatures of a hidden AGN (Marin et al. 2024). It’s also worth noting that the polarization angle is constant in these two objects over time. Again that mitigates the favor of the notion that these two FR Is show their synchrotron cores in the optical. The case is clinched in Centaurus A by the similar high perpendicular polarization in the line complex. In 3C 270, we also see that the complex is polarized at the same polarization angle, though direct line emission must dilute the scattered emission substantial in the latter case since the polarization is only 0.1%.
In Figure 2 of Capetti et al. (2007), the polarization degree is shown as a function of aperture radius to investigate the concentration of polarization. The polarization degree of 3C 270 can reach as high as at the central pixel, consistent with a point source origin. Their polarization degree profile extends to from the center, where it drops to . Given the aperture size in our observation, this result is broadly consistent with our measured continuum polarization.
A key objective of our observations is to determine whether broad H emission, coupled with the Big Blue Bump continuum, can be detected in scattered (polarized) light, as predicted by the unification model. We confirm the detection of polarized emission lines (the H complex) with a polarization angle perpendicular to the jet axis by Barth et al. (1999a). The alignment of both the emission lines and the continuum polarization perpendicular to the jet in our data strengthen a scattering scenario, similar to what has been observed in Centaurus A. We now examine the compositions of the emission line complex in the polarized flux spectrum in detail.
| Wavelength(Å) | (∘) | ||
| 4440 - 5105 | |||
| 5105 - 5770 | |||
| 5770 - 6505 | |||
| 6505 - 6620 (H + [N II]) | |||
| 6620 - 6700 | |||
| 6700 - 6750 ([S II]) | |||
| 6750 - 7440 | |||
| 7440 - 8105 | |||
| 8105 - 9260 |
3.1 Modeling the polarized emission lines with narrow lines only
We first examine if the scattered narrow line emissions solely can account for the features in polarized flux spectrum. To this end, we model the Stokes flux profiles of the H and [S II] complexes using the exact profiles extracted from the total flux spectrum, varying only the normalization. As shown in Fig.4, modeling the polarized line profiles with only narrow lines provides an acceptable fit as far as feature width is concerned. Importantly, no broad component is required in the Stokes flux spectrum to account for underlying emission from the BLR. We acknowledge that the profile in the Stokes flux and total flux differs since the scattered narrow lines see different parts of the mirror when the mirror is comparable with the NLR. However, only non-unique, ad hoc explanations can be given for this peculiar detail.
3.2 Precluding Seyfert I type line profile in the polarized flux spectrum
If the polarized flux contains a Seyfert I-type line profile, can this possibility be robustly ruled out? We adopt the method outlined in Marin et al. (2024) to evaluate what fraction of the genuine type-I profiles would be conspicuous if present in our polarized flux spectrum.
Accurate determination of the polarized continuum flux is essential since the polarization of the continuum is heavily diluted. Using HST imaging polarimetry, where significant polarization has been detected (Capetti et al., 2007), we calculate the polarized flux as a function of aperture radius. The polarized flux flattens at large apertures, yielding an average polarized flux of erg s-1cm-2Å-1. Given that polarized emission lines contribute approximately to this flux, the polarized continuum flux is erg s-1cm-2Å-1, in excellent agreement with our observations.
We scale the real type-I profiles onto the polarized continuum based on their equivalent widths and resample the coadded spectra according to the uncertainties in our polarized flux. The type-I AGN sample is drawn from Schmidt & Green (1983) (23 sources), Stirpe (1990) (5 sources), and Jiang et al. (in prep.; 9 sources). The properties of these sources are summarized in Table A1. This sample comprises highly luminous objects ( mag or log(/erg s-1cm), whose spectra can be compared directly to our polarized flux spectrum due to the absence of host stellar contamination (Shen et al., 2011). We assume that this sample is representative of classified type-I AGNs. We visually inspect the data and find that fewer than 6% (2 out of 37, see Fig.A1) of the type-I profiles would remain undetected.
Therefore, we conclude that the presence of scattered narrow lines is more plausible in 3C 270, and neither a BLR nor a BBB is required under these circumstances.
4 Discussion
Chiaberge et al. (1999) study a sample of FR I radio galaxies using the HST and identify an unresolved optical Compact Central Core (CCC) with a high detection rate, including 3C 270. The presence of a CCC suggests that there is only mild extinction along the line of sight, as substantial extinction would completely obscure the source. Typically, low extinction corresponds to a low column density and further examination on the X-ray emission shows that for most of the sources in the sample, is indeed very low ( cm-2, Balmaverde et al., 2006). However, X-ray studies of 3C 270 require modeling the spectrum with a significantly higher column density of cm-2 (Zezas et al., 2005; Worrall et al., 2010).
The observed discrepancy between high column density and low extinction in 3C 270 could be explained if part of the X-ray obscuring material is not along the line of sight for the optical emission. However, in our favored scenario, we are observing the optical emission through a mirror and do not expect X-ray column density to apply.
5 Conclusion
If the high X-ray column density does mean the optical core is hidden in this object, our data indicate that we can see the hidden nuclear region with a polar mirror. Since we don’t observe broad lines in polarized flux but we do observe continuum and narrow lines, it seems likely that there is no hidden BBB and BLR. Thus, our periscopic view reveals a continuum which is likely beamed synchrotron, and the innermost NLR. The presence of a mid-infrared point source (Asmus et al., 2011, 2014) and warm dust emission (Leipski et al., 2009) in 3C 270 likely arises from the jet or Advection-Dominated Accretion Flow (ADAF) heating the compact torus instead of a BBB. This is our favored hypothesis, making 3C 270 similar to radio-loud AGN Centaurus A (Marin et al. 2024) and radio-quiet AGN NGC 4258 (Barth et al., 1999b). The caveat is that the scattered continuum is only weakly detected, so a commensurate BBB and BLR can’t be totally ruled out.
| Quasar ID | Redshift | Selection Criteria | Reference |
| PG 0157+001 | 0.164 | = -24.83 | 1, 3 |
| PG 0923+201 | 0.190 | = -24.32 | 1, 3 |
| PG 0947+396 | 0.206 | = -24.15 | 1, 3 |
| PG 1001+054 | 0.161 | = -23.86 | 1, 3 |
| PG 1004+130 | 0.240 | = -24.96 | 1, 3 |
| PG 1012+008 | 0.185 | = -24.41 | 1, 3 |
| PG 1048+342 | 0.167 | = -24.26 | 1, 3 |
| PG 1114+445 | 0.144 | = -23.69 | 1, 3 |
| PG 1115+407 | 0.154 | = -23.87 | 1, 3 |
| PG 1116+215 | 0.177 | = -25.03 | 1, 3 |
| PG 1202+281 | 0.165 | = -25.03 | 1, 3 |
| PG 1229+204 | 0.064 | = -23.30 | 1, 3 |
| PG 1307+085 | 0.155 | = -24.63 | 1, 3 |
| PG 1309+355 | 0.184 | = -24.84 | 1, 3 |
| PG 1352+183 | 0.158 | = -24.24 | 1, 3 |
| PG 1404+226 | 0.098 | = -23.07 | 1, 3 |
| PG 1415+451 | 0.114 | = -23.48 | 1, 3 |
| PG 1427+480 | 0.221 | = -24.37 | 1, 3 |
| PG 1519+226 | 0.137 | = -23.54 | 1, 3 |
| PG 1545+210 | 0.266 | = -25.08 | 1, 3 |
| PG 1552+085 | 0.119 | = -23.30 | 1, 3 |
| PG 1612+261 | 0.131 | = -23.53 | 1, 3 |
| PG 2233+134 | 0.326 | = -25.54 | 1, 3 |
| PKS 1020-103 | 0.197 | = -23.53 | 2 |
| Fairall 9 | 0.046 | = -23.25 | 2 |
| PG 0837-12 | 0.198 | = -24.13 | 2 |
| 4C 31.63 | 0.295 | = -25.31 | 2 |
| LEDA 61965 | 0.123 | = -23.30 | 2 |
| SDSS J014017.06-005003.0 | 0.334 | log(/erg s-1) = 45.19 | 3 |
| SDSS J023922.87-000119.6 | 0.262 | log(/erg s-1) = 45.12 | 3 |
| SDSS J082758.52+233302.9 | 0.302 | log(/erg s-1) = 45.12 | 3 |
| SDSS J082958.79+513044.5 | 0.342 | log(/erg s-1) = 45.17 | 3 |
| SDSS J083535.80+245940.1 | 0.330 | log(/erg s-1) = 45.23 | 3 |
| SDSS J100402.60+285535.3 | 0.329 | log(/erg s-1) = 45.30 | 3 |
| SDSS J143204.60+394439.0 | 0.348 | log(/erg s-1) = 45.09 | 3 |
| SDSS J151535.25+480530.5 | 0.312 | log(/erg s-1) = 45.05 | 3 |
| SDSS J161849.24+442517.2 | 0.335 | log(/erg s-1) = 45.03 | 3 |