### v3

# The central engine of low-luminosity radio galaxy 3C 270 (NGC 4261)

Abstract
We present the polarization spectra of the nucleus of 3C 270. We confirm that the polarization angle of both the continuum and the emission lines are close to perpendicular to the jet direction after careful correction of interstellar polarization, which indicates polar scattering. The Stokes flux spectrum resembles the total flux spectrum, with no need for a broad component from the Broad Line Region. Over 94% of a sample of Seyfert I broad line profile would be significantly detected if present in our polarized flux spectrum. We favor the hypothesis that we are observing the continuum including any contribution from Radiatively Inefficient Accretion Flow and beamed synchrotron instead of a Big Blue Bump, as well as the innermost Narrow Line Region, through reflection. This makes 3C 270 the third known case, after NGC 4258 and Centaurus A, where only narrow lines (and the continuum, if present) are scattered, with no evidence of an underlying Big Blue Bump.
1 Introduction
The nearby elliptical galaxy NGC 4261 () hosts a low-luminosity Active Galactic Nuclei (AGN) and also a bright radio source known as 3C 270. This galaxy is renowned for its prominent geometrically thick dust disc (Jaffe et al., 1993) rotating around a supermassive black hole with a mass of , as reported by Boizelle et al. (2021) and Ruffa et al. (2023).
This famous FR I type radio galaxy has been the subject of numerous studies trying to establish its accretion mode : radiatively efficient, ”thermal” vs. radiatively inefficient, ”non-thermal” (Antonucci, 2012). This interest is motivated by the fact that 3C 270 appears to be a low-luminosity active galaxy showing puzzling evidence for significant accretion power, making it an ideal testing ground for our current understanding of accretion onto supermassive black holes at low Eddington ratios. On one hand, the low-efficiency accretion is supported by emission-line diagnostics from ground-based spectra, which classify 3C 270 as a low-ionization nuclear emission-line region (LINER) (Ho et al., 1995; Jaffe et al., 1996). However, Hubble Space Telescope (HST) spectra show Seyfert-type narrow-line ratios on nuclear scales (Ferrarese et al., 1996), indicating the presence of a hard radiation field typical for radiatively efficient accretion 111The nucleus predominantly exhibits a LINER spectrum, but space-based observations within a central 0.1′′ aperture reveal nuclear line ratios ([O III] / H 5) that fall in the Seyfert region on the [S II] and [N II] BPT diagrams (Kewley et al., 2006; Kauffmann et al., 2003), suggesting the possible presence of a hard radiation field. Old stellar population (Wills et al., 2004) with a low star formation rate ( 0.01 M⊙ yr-1, Vaddi et al., 2016) indicates that underlying Balmer absorption from host galaxy ( 1.5Å, Worthey, 1994) does not significantly affect the line ratios (Equivalent Width of H 10Å) on the BPT diagram..
From Spitzer spectroscopy, Leipski et al. (2009) showed that 3C 270 is dominated by warm thermal dust ( 240 K) in the mid-infrared, too warm for stellar-heated dust, and indicative of a possible hidden thermal AGN (Ogle et al., 2006). Very Large Telescope Imager and Spectrometer for mid Infrared (VLT-VISIR) data at 12 micron reveal a matching point source, which proves that the warm thermal dust has indeed a nuclear origin ( 60 pc; Asmus et al., 2011). However, in some clumpy torus models, a powerful nuclear thermal dust component is not expected for sources of low luminosity. Instead, the models predict that the torus (and the broad line region) vanishes at low bolometric luminosity ( 1042 erg s-1) with 3C 270 having erg s-1 (Ho, 1999) and/or at low Eddington ratios (Hönig & Beckert, 2007; Elitzur, 2008; Elitzur & Ho, 2009; González-Martín et al., 2017). Additionally, for the very low Eddington ratio of 3C 270 (, Ho, 1999), theory predicts advection-dominated accretion (Esin et al., 1997), thus limiting the supply of hard ionizing radiation. However, the nuclear optical emission lines in 3C 270 requires a central engine to produce sufficient ionizing photons. Furthermore, the reprocessed light observed as nuclear warm dust emission necessitates a radiation field sufficient to power the emission. These discrepancies illustrate the need to establish more accurate observational constraints to the models.
One important diagnostic that could place strong constraints on the models for low-luminosity activity is the detection of broad emission lines and the Big Blue Bump continuum component. These are generally considered reliable probes for the presence of a radiatively efficient accretion flow, and the detection of broad lines in polarized light would be the most convincing proof for a hidden thermal AGN. Spectropolarimetry has been used successfully many times to reveal hidden broad lines in narrow-line objects (e.g., Antonucci, 1984; Ogle et al., 1997; Ramos Almeida et al., 2016). Additionally, 3C 270 has been observed spectropolarimetrically by Barth et al. (1999a), surprisingly revealing possibly broadened H in polarized light. Unfortunately, the low signal-to-noise ratio (SNR) precluded a secure detection and width measurement of the apparent emission line feature. In concert with the detection of warm nuclear dust emission, establishing the presence of broad emission lines in 3C 270 would provide strong constraints for our current models of low-luminosity activity, which predict both of these features to be absent. It should be noted however that the theoretical expectations are already known to be contradicted in some cases (e.g., NGC 1052, Barth et al., 1999b).
In this work, we present a deeper optical spectropolarimetric observation on 3C 270, aiming to disentangle the H complex in search of hidden broad line. In Section 2, we describe the observations and data reduction and careful removal of interstellar polarization. In Section 3, we analyze different models for the polarized flux spectrum. In Section 4, we discuss the implication of accretion mode as from the polarized flux spectrum and make a summary. Throughout the paper, we adopt a CDM cosmology with = 0.7, = 0.3, and H0 = 70 km s-1 Mpc-1.
2 Observations and Data Reduction
The observations of 3C 270 were conducted using the FOcal Reducer and low dispersion Spectrograph (FORS2), mounted on the Unit Telescope 1 of the Very Large Telescope (VLT) at the European Southern Observatory (ESO). The observations were spread over three nights: 2014 / 01 / 24, 2014 / 02 / 28, and 2014 / 03 /01 (Program ID: 092.B-0401A). The seeing conditions vary within the three nights as follows: 0.9′′ - 1.1′′ (2014 / 01 / 24); 1.2′′ - 1.7′′ (2014 / 02 / 28); 0.5′′ - 0.9 ′′(2014 / 03 / 01). All data were taken in PMOS mode with the 300V grism.
The observations employed the Half-Wave Plate (HWP) at four different position angles of 0.0∘, 22.5∘, 45.0∘, and 67.5∘ 222The HWP angles correspond to 0.0∘, 45.0∘, 90.0∘, and 135.0∘ on the sky plane.. A Wollaston prism was used to split the light into ordinary ray (o-ray) and extraordinary ray (e-ray). Due to the dual-beam design of FORS2, both o-ray and e-ray were recorded simultaneously, with a separation of about 22 arcseconds. The total on-source exposure time for each half-wave plate position was 3000s.
The data were reduced using routines written in IDL, following the procedures described by Martínez González et al. (2015) and Ramos Almeida et al. (2016). The steps included bias subtraction, flat-field correction, wavelength calibration, sky subtraction, cosmic ray removal, and flux calibration. Both o-ray and e-ray spectra are extracted using an aperture with a radius equal to the FWHM of the seeing for each epoch, and then combined with weights based on the exposure time of each epoch. The o-ray and e-ray spectra were then combined to compute the normalized Stokes parameters and using the following formulas:
where is the measured intensity, with the first subscript indicating the ordinary (o) or extraordinary (e) ray, and the second subscript specifying the HWP rotation angle.
The total intensity is computed as: .
The linear polarization degree and polarization angle PA are derived from and using the standard relations:
For calibration, we selected a zero-polarization standard star (HD 109055) and a polarized standard star (Vela1 95), both observed on 2024 / 01 / 24 after the first source observation. Flux and polarimetric standards were part of the default nightly calibration plan.
In the R band, Vela1 95 has a polarization degree of and a polarization angle of 333https://www.eso.org/sci/facilities/paranal/instruments/fors/inst/pola.html. For comparison with the standard, we integrate the normalized Stokes parameters and from our data, weighted by the Johnson R filter transmission. From our measurements, the polarization degree is 7.79 0.04% with a polarization angle of , very close to the standard values. Using the same procedure, we measure the polarization degree of HD 109055, finding 0.06 0.04%, consistent with expected null polarization. The close agreement between our measured and tabulated values indicates that the instrumental polarization in the R band is .
The total flux spectrum as well as the Stokes and spectra before interstellar polarization correction are shown in Fig. 1.
2.1 Interstellar Polarization
Since the continuum polarization is quite low, interstellar polarization (ISP) may significantly impact our interpretation of its origin. Serkowski et al. (1975) demonstrate that the wavelength-dependent polarization degree of the ISP can be accurately modeled using two parameters, and . With the additional constraint of the ISP polarization angle, we can derive the complete ISP information, including the normalized (fractional) Stokes parameters and . Key parameters for the ISP can be estimated from polarization measurements of stars close to the central source, which provides a local basis for ISP correction.
We start with the polarization catalog of Galactic stars by Heiles (2000) and take the polarization properties of the closest star to the source to serve as a proxy for the ISP of the source. We also require the proxy star to be sufficiently distant from the Earth to maximally sample the ISP caused by Galactic dust along the line of sight. The best candidate matching these criteria is HD 107289, located 300 pc away from the Earth at Galactic latitude of and 12′ away from 3C 270 in projection. Galactic ISP exhibits intrinsic asymmetry on large scale (Berdyugin et al., 2014). However, their Fig. 3 shows that the polarization angle of ISP is largely coherent around 3C 270. We verify this coherence by comparing the polarization angle of HD 107289 to that of the nearby stars within 5∘, finding variations within 10∘, which will not affect our analysis. Therefore, HD 107289 is a good proxy for measuring Galactic ISP. The measured polarization degree is 0.084% 0.012% and the polarization angle is . To check for residual dust between HD 107289 and 3C 270 , we use the 3D dust map from Green et al. (2019) to calculate the integrated extinction along the line of sight. In the direction of 3C 270, the integrated Galactic extinction is = 0.059 0.014 mag, while toward HD 107289, the value is 0.047 0.018 mag. This indicates that there is minimal dust beyond HD 107289 that would contribute to ISP within our Galaxy. Taking = 0.084%, = 64∘ and assuming is 5500 Å, we can then construct the synthetic normalized and spectra as follows:
The computed normalized Stokes parameters are then subtracted from the observed normalized and to derive the intrinsic values. The error of the intrinsic normalized Stokes parameters is calculated by combining the ISP error and the measurement error in quadrature. The normalized Stokes parameter and are shown in Fig. 2.
3 Results
The radio jet position angle is measured at 4.885 GHz (Birkinshaw & Davies, 1985). After correcting for the ISP, we find that the polarization angle of the continuum is close to perpendicular to the jet direction, as shown in Fig. 3 (see also Table 1). At wavelength of 6000 Å, we found a continuum polarization angle of , aligning well with the findings of Capetti et al. (2007) at matching wavelength. Their broadband polarimetry, primarily tracing continuum polarization, shows a polarization angle of , perpendicular to the jet. Most importantly, the polarization angle of both the H and [S II] complexes, which are less subject to ISP correction, are and , very closely perpendicular to the jet axis at 88∘. Our results are in excellent agreement with theirs, demonstrating the precision of our ISP correction as well.
Capetti et al. (2007) report the optical polarization of a sample of FR I radio galaxies. In particular, their Table 2 shows an overall uncorrelated polarization angles to the jet position angle. The lack of consistent perpendicularity across their sample leads Capetti et al. (2007) to attribute the observed polarization to intrinsic radiation mechanisms, most likely synchrotron emission. However, a notable distinction between 3C 270 and the other FR Is in Capetti et al. (2007), besides its perpendicular polarization, is the high X-ray column density of cm-2 (Zezas et al., 2005; Worrall et al., 2010). Thus, 3C 270 is similar to the FR I prototype Centaurus A, which wasn’t observed by Capetti et al. (2007), but which also has a perpendicular polarization and a high column density associated with the nuclear region, two signatures of a hidden AGN (Marin et al. 2024). It’s also worth noting that the polarization angle is constant in these two objects over time. Again this argues against the notion that these two FR Is show their synchrotron cores directly in the optical. The case is clinched in Centaurus A by the similar high perpendicular polarization in the line complex. In 3C 270, we also see that both H and [S II] complex are polarized at the same polarization angle, though direct line emission must dilute the scattered emission substantially in the latter case since the polarization is only 0.1%.
In Fig. 2 of Capetti et al. (2007), the polarization degree is plotted as a function of aperture radius to investigate the concentration of polarization. The polarization degree of 3C 270 reaches as high as 10% at the central pixel, consistent with a point source origin, and drops to 0.2% at an aperture radius of . Given that our observation uses a comparable aperture size to the latter, this polarization degree is broadly consistent with our measured continuum polarization.
A key objective of our observations is to determine whether broad H emission, coupled with the Big Blue Bump continuum, can be detected in scattered (polarized) light, as predicted by the unification model. We confirm the detection of polarized narrow emission lines (the H complex) with a polarization angle perpendicular to the jet axis by Barth et al. (1999a). The alignment of both the emission lines and the continuum polarization perpendicular to the jet in our data strengthen a scattering scenario, similar to what has been observed in Centaurus A. We now examine the compositions of the emission line complex in the polarized flux spectrum in detail.
| Wavelength(Å) | PA (∘) | ||
| 4440 - 5105 | |||
| 5105 - 5770 | |||
| 5770 - 6505 | |||
| 6505 - 6620 (H + [N II]) | |||
| 6620 - 6700 | |||
| 6700 - 6750 ([S II]) | |||
| 6750 - 7440 | |||
| 7440 - 8105 | |||
| 8105 - 9260 |
3.1 Modeling the polarized emission lines with narrow lines only
We first examine if the scattered narrow line emissions solely can account for the features in polarized flux spectrum. To this end, we model the Stokes flux profiles of the H and [S II] complexes using the exact profiles extracted from the total flux spectrum, varying only the normalization. As shown in Fig. 4, modeling the polarized line profiles with only narrow lines provides an acceptable fit ( as far as feature width is concerned. Importantly, no broad component is required in the Stokes flux spectrum to account for underlying emission from the BLR. We acknowledge that the profile in the Stokes flux and total flux may differ in detail since the scattered narrow lines see different parts of the mirror when the size of the mirror is comparable with the NLR. However, only non-unique, ad hoc explanations can be given for this peculiar detail.
Traditional methods that place an upper limit on the broad H flux from the BLR require assuming the profile of the undetected line, which can lead to erroneous constraints (e.g., NGC 3147, Bianchi et al. 2019, 2022; see also the discussion in Antonucci 2012). Therefore, in the next section, we adopt a different approach, aiming not to rigorously constrain the H contribution from the BLR, but to robustly rule out the presence profiles typical of Seyfert I AGN.
3.2 Precluding Seyfert I type line profile in the polarized flux spectrum
Is it possible to robustly rule out the possibility that the polarized flux contains a Seyfert I-type line profile? We adopt the method outlined in Marin et al. (2025) to evaluate what fraction of the genuine type-I profiles would be conspicuous if present in our polarized flux spectrum.
Accurate determination of the polarized continuum flux is essential since the polarization of the continuum is heavily diluted by starlight in total flux. Using HST imaging polarimetry, where significant polarization has been detected (Capetti et al., 2007), we obtain a polarized flux of erg s-1 cm-2 Å-1 within 1.2′′ radius aperture, comparable to the average seeing of our observations. Given that polarized narrow emission lines contribute approximately 20% to this flux with the same polarization angle, the polarized continuum flux is erg s-1 cm-2 Å-1, in excellent agreement with our observations.
We scale the real type-I profiles onto the polarized continuum based on their equivalent widths and resample the coadded spectra according to the uncertainties in our polarized flux. The type-I AGN sample is drawn from Schmidt & Green (1983) (23 sources), Stirpe (1990) (5 sources), and Jiang et al. (in prep.; 9 sources). The properties of these sources are summarized in Table A1. This sample comprises highly luminous objects ( mag or log(/erg s-1 cm), whose spectra can be compared directly to our polarized flux spectrum due to the absence of host stellar contamination (Shen et al., 2011). We assume that this sample is representative of classified type-I AGNs. We calculate the SNR of the broad wing region in the coadded spectra, defined at the wavelength corresponding to the half-maximum. A broad feature is considered detected if the SNR at this wavelength exceeds 3. We find that fewer than 6% (2 out of 37, see Fig. A1) of the type-I profiles would remain undetected.
Therefore, we conclude that the presence of scattered narrow lines is more plausible in 3C 270, and neither a BLR nor a BBB is required under these circumstances.
4 Discussion and Summary
Chiaberge et al. (1999) studied a sample of FR I radio galaxies using the HST and identified an unresolved optical Compact Central Core (CCC) with a high detection rate, including 3C 270. The presence of a CCC in 3C 270 suggests that there is only mild extinction ( 3) along the line of sight, as substantial extinction would completely obscure the source (Chiaberge et al., 2003). Typically, low extinction corresponds to a low column density and further examination on the X-ray emission shows that for most of the sources in the sample, is indeed very low ( cm-2, Balmaverde et al., 2006). However, X-ray studies of 3C 270 require modeling the spectrum with a significantly higher column density of cm-2 (Zezas et al., 2005; Worrall et al., 2010).
The observed discrepancy between high column density and low extinction in 3C 270 could be explained if most of the X-ray obscuring material is not along the line of sight to the optical emission. In our favored scenario, we are observing the optical emission through a mirror and do not expect X-ray column density to apply.
If the high X-ray column density does mean the optical core is hidden in this object, our data indicate that we can see the hidden nuclear region with a polar mirror. Since we don’t observe broad lines in polarized flux but we do observe continuum and narrow lines, it seems likely that there is no hidden BBB and BLR. Thus, our periscopic view reveals the underlying continuum, including any contributions from Radiatively Inefficient Accretion Flow (RIAF) and beamed synchrotron, and the innermost NLR. With the available data, we can not distinguish which of the two components dominates the scattered continuum. In the simplest FR I – BL Lac unification scheme (Chiaberge et al., 1999; Capetti et al., 2000; Trussoni et al., 2003), a scattered continuum dominated by beamed synchrotron emission would be favored. de Menezes et al. (2020) attempted a broadband SED analysis of 3C 270 and favored a jet-dominated synchrotron self - Compton model over the RIAF model. However, we caution that the optical emission being scattered should be fully considered when performing such analyses, and that multi-wavelength (polarimetric) SED analysis are needed to better distinguish the two components (e.g., Marin, 2018). In either cases, the presence of a mid-infrared point source (Asmus et al., 2011, 2014) and warm dust emission (Leipski et al., 2009) in 3C 270 arising from the BBB heating the compact torus can be ruled out. This is our favored hypothesis, making 3C 270 similar to radio-loud AGN Centaurus A (Marin et al., 2025) and radio-quiet AGN NGC 4258 (Barth et al., 1999c). The caveat is that the scattered continuum is only weakly detected, so a commensurate BBB and BLR can’t be totally ruled out.
| Quasar ID | Redshift | Selection Criteria | Reference |
| PG 0157+001 | 0.164 | = -24.83 | 1, 3 |
| PG 0923+201 | 0.190 | = -24.32 | 1, 3 |
| PG 0947+396 | 0.206 | = -24.15 | 1, 3 |
| PG 1001+054 | 0.161 | = -23.86 | 1, 3 |
| PG 1004+130 | 0.240 | = -24.96 | 1, 3 |
| PG 1012+008 | 0.185 | = -24.41 | 1, 3 |
| PG 1048+342 | 0.167 | = -24.26 | 1, 3 |
| PG 1114+445 | 0.144 | = -23.69 | 1, 3 |
| PG 1115+407 | 0.154 | = -23.87 | 1, 3 |
| PG 1116+215 | 0.177 | = -25.03 | 1, 3 |
| PG 1202+281 | 0.165 | = -25.03 | 1, 3 |
| PG 1229+204 | 0.064 | = -23.30 | 1, 3 |
| PG 1307+085 | 0.155 | = -24.63 | 1, 3 |
| PG 1309+355 | 0.184 | = -24.84 | 1, 3 |
| PG 1352+183 | 0.158 | = -24.24 | 1, 3 |
| PG 1404+226 | 0.098 | = -23.07 | 1, 3 |
| PG 1415+451 | 0.114 | = -23.48 | 1, 3 |
| PG 1427+480 | 0.221 | = -24.37 | 1, 3 |
| PG 1519+226 | 0.137 | = -23.54 | 1, 3 |
| PG 1545+210 | 0.266 | = -25.08 | 1, 3 |
| PG 1552+085 | 0.119 | = -23.30 | 1, 3 |
| PG 1612+261 | 0.131 | = -23.53 | 1, 3 |
| PG 2233+134 | 0.326 | = -25.54 | 1, 3 |
| PKS 1020-103 | 0.197 | = -23.53 | 2 |
| Fairall 9 | 0.046 | = -23.25 | 2 |
| PG 0837-12 | 0.198 | = -24.13 | 2 |
| 4C 31.63 | 0.295 | = -25.31 | 2 |
| LEDA 61965 | 0.123 | = -23.30 | 2 |
| SDSS J014017.06-005003.0 | 0.334 | log(/erg s-1) = 45.19 | 3 |
| SDSS J023922.87-000119.6 | 0.262 | log(/erg s-1) = 45.12 | 3 |
| SDSS J082758.52+233302.9 | 0.302 | log(/erg s-1) = 45.12 | 3 |
| SDSS J082958.79+513044.5 | 0.342 | log(/erg s-1) = 45.17 | 3 |
| SDSS J083535.80+245940.1 | 0.330 | log(/erg s-1) = 45.23 | 3 |
| SDSS J100402.60+285535.3 | 0.329 | log(/erg s-1) = 45.30 | 3 |
| SDSS J143204.60+394439.0 | 0.348 | log(/erg s-1) = 45.09 | 3 |
| SDSS J151535.25+480530.5 | 0.312 | log(/erg s-1) = 45.05 | 3 |
| SDSS J161849.24+442517.2 | 0.335 | log(/erg s-1) = 45.03 | 3 |